# Learning from Emergence: A Study on Proactively Inhibiting the Monosemantic Neurons of Artificial Neural Networks

### v2

@@ Abstract.
Recently, emergence has received widespread attention from the research community along with the success of large language models. Different from the literature, we hypothesize a key factor that highly promotes the performance during the increase of scale: the reduction of monosemantic neurons that can only form one-to-one correlations with specific features. Monosemantic neurons tend to be sparser and have negative impacts on the performance in large models. Inspired by this insight, we propose an intuitive idea to identify monosemantic neurons and inhibit them. However, achieving this goal is a non-trivial task as there is no unified quantitative evaluation metric and simply banning monosemantic neurons does not promote polysemanticity in neural networks. Therefore, we propose to learn from emergence and present a study on proactively inhibiting the monosemantic neurons in this paper. More specifically, we first propose a new metric to measure the monosemanticity of neurons with the guarantee of efficiency for online computation, then introduce a theoretically supported method to suppress monosemantic neurons and proactively promote the ratios of polysemantic neurons in training neural networks. We validate our conjecture that monosemanticity brings about performance change at different model scales on a variety of neural networks and benchmark datasets in different areas, including language, image, and physics simulation tasks. Further experiments validate our analysis and theory regarding the inhibition of monosemanticity.
Deep Learning, Artificial Neural Networks, Emergent Ability, Monosemanticity
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; August, 2024; Spainisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Figure 1. Demonstration of important concepts:(a) A multilayer neuron network for the image classification task. The middle layer is highlighted. (b) A monosemantic neuron (orange) ideally activates for one specific type of feature. (c) A polysemantic neuron activates for multiple features. (d) A distributed feature is classified based on the activation of multiple neurons.
@@
The activity of artificial neural networks diminished for decades before experiencing great success after the 2010s (Krogh, 2008; Krizhevsky et al., 2012). One major difference compared to previous models is the increasing scale. In recent years, large neural networks have much larger scales in terms of datasets, model sizes, and training quantities, which have achieved remarkable results in various fields (Floridi and Chiriatti, 2020; Touvron et al., 2023). Given increasing scales, the performance improvements can usually be estimated under the scaling law (Kaplan et al., 2020), which follows stable but slow increasing trend of power laws (OpenAI, 2023). However, improvements that significantly deviate from estimation occur when the scale increases. Emergence, an intriguing phenomenon in large-scale models, refers to the gradual improvement of model performance before the scale reaching a certain threshold, followed by a rapid enhancement once the threshold is surpassed (Wei et al., 2022). Increasing evidence suggests that the surprises may not arise from new module and architecture designs, but rather from the underlying nature of scale changes (Zhou et al., 2023). A critical question arises: people increase the model scale and get better results, but what has changed underlying the process?
@@
From the perspectives of monosemantic and polysemantic neurons, some pioneer works try to interpret the performance of small and large-scale models, respectively. Through statistical analysis of the relationships between neurons and input features, a neuron is considered monosemantic if it forms a 1-1 correlation with its related input feature (Bau et al., 2020; Elhage et al., 2022) (Figure 1-b). In contrast, polysemantic neurons activate for several features that are weakly correlated (Trenton et al., 2023; Chris et al., 2020; Gabriel et al., 2021; Elhage et al., 2022) (Figure. 1-c). Researchers find that smaller models incur larger errors (Gurnee et al., 2023) when an input’s related monosemantic neuron is disabled. Following the favor of monosemantic neurons from studies on small-scale models, existing works focus on enhancing or extracting monosemanticity (Gurnee et al., 2023; Elhage et al., 2022; Trenton et al., 2023). However, some evidences support that monosemantic neurons are sparser in larger models (Gurnee et al., 2023; Trenton et al., 2023). Those observations imply a positive relationship between the decrease in monosemanticity and the increase in model scale. In other words, monosemanticity could be negatively related to large-scale and good performance. In this paper, inspired by those observations, we propose an important hypothesis that the decrease of monosemantic neurons is a key factor in achieving higher performance as the model scale increases. Figure. 4 in Appx. A shows an example analogy to demonstrate our idea.
@@ 2. Preliminary
where $\mathbf{x}=[x_{1},\cdots,x_{d}]\in\mathbb{R}^{d}$ is a $d$-dimensional vector representing the $n$-th input data sample, such as a weather record series covering $d$ days or an image with height $\times$ width $=d$ pixels.
Without loss of generality, let $f(\mathbf{W};\mathbf{x})$ be a $L$-th layer fully connected neural networks (Haykin, 1994). We denote $h_{i}^{\ell}$, $\sigma_{i}^{\ell}(\cdot)$, $z_{i}^{\ell}$ as the input after linear transformation, activation function, and output of the $i$-th hidden unit at $\ell$-th layer, respectively. Then, we may formally define the computation within neurons as follows:
@@ 2.1. Activation and Monosemantic
$\forall_{\mathbf{x}}act(z_{i},\mathbf{x})=1,\mathbf{x}\in C_{j}$
However, features are human-defined and vary a lot. In their study, Gurnee et al. (2023) consider about 100 features, while Trenton et al. (2023) study up to $10^{5}$ features to fully capture monosemanticity. Without unified feature datasets, it is also difficult to explicitly formalize the definition of “monosemantic”. More related works are discussed in Appx. B.
@@ 3.1. Metric for Monosemanticity
The proof is given in Appendix. C.1. In the implementation, we introduce a forgetting mechanism during model updates, where the influence of previous samples should decay. For details, please refer to Algorithm. 1 in Appendix. E.2.
@@ 3.2.1. Naive Deactivation
Recall that given a monosemantic neuron $z$ and an input $\mathbf{x}$ that contains its exclusive feature, we aim to optimize the model in two aspects: (1) decrease the activation level of $z$ when given $\mathbf{x}$ (Figure. 2(b)); and (2) reduce the reliance of output $\mathbf{o}$ on the activation of $z$ (Figure. 2(c)).
@@
where subscript $\cdot_{ng}$ refers to a “no gradient” fixed-value scalar without trainable parameters, compared with $z\in\mathbf{z}=f_{1}(\mathbf{x})$ which is updatable. Two examples are given in Figure. 2(e,f). Both solutions ensure that $f_{2}$ receives an updated $z^{\prime}$, in which the value of $\mathcal{V}(z^{\prime})=\bar{z}$ provides little information. As $z^{\prime}$ is valueless, $f_{2}$ has to adjust its parameter to utilize information from other neurons in $\mathbf{z}$ for good output $\mathbf{o}^{*}=f_{2}^{*}(\mathbf{z}^{\prime},\mathbf{x})$. Such a process achieves our second goal of reducing the dependence of output $\mathbf{o}$ on the activation of $z$ (Figure. 2(l)).
However, neither of the two solutions can achieve the first goal (i.e., training $f_{1}$ to inhibit the activation of $z$ for a single feature). To be more specific, method (a) wastes all the gradient for $f_{1}$ in generating $z$, preventing the related parameters from being updated. Method (b) outputs $\mathcal{V}(z^{\prime})=\bar{z}$ by compensating for the gap. As the model obtains a good result with $z$, by calculating the gradient, $f_{1}$ will be updated to push its output value from $\bar{z}$ to $z$, expressed as $\bar{z}+\delta(z-\bar{z})$, where $\delta>0$ up to the learning rate (Figure. 2(i)). Ironically, the actual output of $z$ is updated to $z+\delta(z-\bar{z})$, deviating from $\bar{z}$ by a factor of $1+\delta$:
@@
Therefore, these two simple solutions either contribute nothing to the deactivation of $z$ or even enhance its activation (Figure. 2(gk)).
@@ 3.2.2. Reversed Deactivation
To address the aforementioned problems, we propose our method called Reversed Deactivation (RD for short). Following the above-mentioned definitions, we replace the original $z$ with $z^{\prime}$ (Figure. 2(f)):
@@
Similar to baselines, RD also ensures the second goal of decreasing the dependence of output $\mathbf{o}$ on the activation of $z$. To be more specific, $f_{2}$ receives a value $\mathcal{V}(z^{\prime})=\mathcal{V}({-z+(\overline{z}+z)_{ng}})=\overline{z}$, which provides little information and requires $f_{2}$ to learn the information from other neurons in $\mathbf{z}$ (Figure. 2(h)).
Besides, RD can inhibit the activation level of $z$ when given $\mathbf{x}$. In short, after calculating the gradient, $f_{1}$ will update the trainable parameter to push its output value from $\bar{z}$ to $z$ (Figure. 2(i)). Different from method (b), the gradient on $z$ is reversed. To achieve the same shifted value $\bar{z}+\delta(z-\bar{z})$ mentioned above, an insight into the update can be expressed as:
@@
which achieves deactivation (Figure. 2(j)). The formal and detailed theory is presented in the following lemma.
Lemma 3.3 ().
Given a trained model $f$ with $2$ continuous derivatives and a Lipschitz continuous gradient, where $f$ achieves a desired output $\mathbf{o}$ with minimal loss $\mathcal{L}(\mathbf{o})$, in which $\textbf{o}=f(\mathbf{x})=f_{2}(f_{1}(\mathbf{x}),\mathbf{x})=f_{2}(\mathbf{z},\mathbf{x})$ for input $\mathbf{x}$ based on its monosemantic neuron $z$ in layer $\mathbf{z}$, suppose that $\mathcal{L}(f_{2}(\cdot))$ monotonically increases with $|z^{\prime}-z|$ for any other value $z^{\prime}$ that replaces $z$. Then, with a sufficiently small learning rate $l$, by updating the model $f$ with gradient descent based on the neuron processed by the RD method, the activation of $z$ on input $\mathbf{x}$ can be inhibited.
The proof is given in Appendix. C.2.
Additionally, we conduct validation experiments to compare the optimization outputs based on different methods in subsection. 4.3. The results are consistent with the theory and the analysis of naive methods and RD, showing that RD is effective in inhibiting monosemanticity.
@@ 3.3. Flexible Plug-in Module
The details of MEmeL are displayed in Figure. 3. We present a general framework of a neural network in Figure. 3(a). The output $\mathbf{o}$ is derived from $\mathbf{x}$ based on the layers of neurons $\{\mathbf{z}\}$. Yellow arrows indicate the reliance of neuron layers on each other. Our method can be applied to any layer of neurons, such as $\mathbf{z}^{3}$ and $\mathbf{z}^{5}$ in Figure. 3(b), where the original neurons are adjusted by our modules to inhibit monosemanticity.
Taking $\mathbf{z}^{5}$ as an example (Figure. 3(c)), we first detect monosemantic neurons using our MS metric, as introduced in subsection. 3.1. After identifying these neurons (colored red), we apply our Reverse Deactivation method, as described in subsection. 3.2, to each of them.
@@
For lightweightness, neither of our two methods introduces any trainable parameters. Additionally, MEmeL focuses on presenting the idea of functionality induction. The methods we propose do not directly decrease monosemanticity during training, but instead induce the model to do so, supported by theoretical analysis. Once we have a well-trained model, the performance is expected to be robust even without induction. Therefore, during testing, the module can be directly removed without incurring any inference overhead. The results of the validation experiments also support our analysis of removing MEmeL during testing. (see Appx. F.1).
@@
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 66.4 80.0
MEmeL-Tune 84.8/83.9 71.7 91.2 93.7 55.7 86.6 89.0 68.1 80.5
The detailed algorithm is provided in the Appendix. D.2. We also apply two simple tricks for implementation: late start and variance compensation, to avoid unstable value fluctuations during the cold start.
@@ 4. Experiments
In this section, we first introduce the basic experimental setup in Sec. 4.1, then present the main empirical studies in Sec. 4.2. We show a case study that the proposed Reverse Deactivation is more powerful in inhibition monosematicity than the naive methods in Sec. 4.3. Furthermore, we discuss the potentials and limitations of MEmeL in Sec. 4.4. Note the experiment about the impact of removing MEmeL during test is provided in Appx. F.1 due to the limited space.

### v3

@@ Abstract.
Recently, emergence has received widespread attention from the research community along with the success of large-scale models. Different from the literature, we hypothesize a key factor that promotes the performance during the increase of scale: the reduction of monosemantic neurons that can only form one-to-one correlations with specific features. Monosemantic neurons tend to be sparser and have negative impacts on the performance in large models. Inspired by this insight, we propose an intuitive idea to identify monosemantic neurons and inhibit them. However, achieving this goal is a non-trivial task as there is no unified quantitative evaluation metric and simply banning monosemantic neurons does not promote polysemanticity in neural networks. Therefore, we first propose a new metric to measure the monosemanticity of neurons with the guarantee of efficiency for online computation, then introduce a theoretically supported method to suppress monosemantic neurons and proactively promote the ratios of polysemantic neurons in training neural networks. We validate our conjecture that monosemanticity brings about performance change at different model scales on a variety of neural networks and benchmark datasets in different areas, including language, image, and physics simulation tasks. Further experiments validate our analysis and theory regarding the inhibition of monosemanticity.
Deep Learning, ANN, Emergent Abilities, Monosemanticity
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671776isbn: 979-8-4007-0490-1/24/08ccs: Theory of computation Machine learning theoryccs: Theory of computation Mathematical optimization
@@ 1. Introduction
Figure 1. Demonstration of important concepts with statistics: (a) A monosemantic neuron (orange) ideally activates for one specific type of feature. (b) A polysemantic neuron (green) activates for multiple features. (c) The output values of a monosemantic neuron when different features are inputted. Its related feature (French) produces values that significantly stand out from other features. (d) The output values of an arbitrarily selected neuron (layer 3, number 333) given different features. The values fluctuate slightly with similar patterns. These statistics are obtained by inspecting the Pythia-v0 410M model (Biderman et al., 2023).
Figure 2. We detect the monosemantic neurons of “French” following the sparse probing paper (Gurnee et al., 2023) and run the experiments on Pythia-v0 (Biderman et al., 2023). After deactivating a monosemantic neuron for “French”, there is an increase in the loss given inputs of different language features (e.g., Dutch and Greek) on Pythia models of different scales: (a) on the 70M Pythia-v0 model, (b) on the 1B Pythia-v0 model, and (c) on the 6.9B Pythia-v0 model. It can be observed that these neurons are typically monosemantic, causing a large increase in loss only when the input contains “French” (see green arrow). However, for larger models, deactivation of these neurons leads to a smaller increase in loss (see red arrow). This gives us a hint that monosemanticity may be negatively related to the scale and performance of larger models.
@@
The activity of artificial neural networks diminished for decades before experiencing great success after the 2010s (Krogh, 2008; Krizhevsky et al., 2012). One major difference compared to previous models is the increasing scale. In recent years, large neural networks have much larger scales in terms of datasets, model sizes, and training quantities, which have achieved remarkable results in various fields (Floridi and Chiriatti, 2020; AI, 2023). Given increasing scales, the performance improvements can usually be estimated under the scaling law (Kaplan et al., 2020), which follows stable but slow increasing trend of power laws (OpenAI, 2023). However, improvements that significantly deviate from estimation occur when the scale increases. Emergence, an intriguing phenomenon in large-scale models, refers to the gradual improvement of model performance before the scale reaching a certain threshold, followed by a rapid enhancement once the threshold is surpassed (Wei et al., 2022). Increasing evidence suggests that the surprises may not arise from new module and architecture designs, but rather from the underlying nature of scale changes (Zhou et al., 2023). A critical question arises: people increase the model scale and get better results, but what has changed underlying the process?
@@
From the perspectives of monosemantic and polysemantic neurons, some pioneer works try to interpret the performance from small to large-scale models. Through statistical analysis of the relationships between neurons and input features, a neuron is considered monosemantic if it forms a 1-1 correlation with its related input feature (Bau et al., 2020; Elhage et al., 2022) (Figure 1(a)). In contrast, polysemantic neurons activate for several features that are weakly correlated (Trenton et al., 2023; Chris et al., 2020; Gabriel et al., 2021; Elhage et al., 2022) (Figure 1(b)). By comparing Figure 1(c) and (d), we can observe that the activation level of a monosemantic neuron for the feature “French” is significantly different from that of the neuron for other features and from that of other non-monosemantic neurons for “French”. Researchers find that smaller models incur larger errors (Gurnee et al., 2023) when an input’s related monosemantic neuron is disabled. Following the favor of monosemantic neurons from studies on small-scale models, existing works focus on enhancing or extracting monosemanticity (Gurnee et al., 2023; Elhage et al., 2022; Trenton et al., 2023). However, some evidences support that monosemantic neurons are sparser in larger models (Gurnee et al., 2023; Trenton et al., 2023). Those observations imply a positive relationship between the decrease in monosemanticity and the increase in model scale. In other words, monosemanticity could be negatively related to large-scale and good performance. In Figure 2, we conduct experiments to show that, when a monosemantic neuron for “French” is deactivated, a smaller model has a greater increase in loss than a larger model. In this paper, inspired by those observations, we propose an important hypothesis that the decrease of monosemantic neurons is a key factor in achieving higher performance as the model scale increases. Figure 5 in Appendix A shows an example analogy to demonstrate our idea.
@@ 2. Preliminary
where $\mathbf{x}=[x_{1},\cdots,x_{d}]\in\mathbb{R}^{d}$ is a $d$-dimensional vector representing the $n$-th input data sample, such as a weather record series covering $d$ days or an image with height $\times$ width $=d$ pixels. Without loss of generality, let $f(\mathbf{W};\mathbf{x})$ be a $L$-th layer fully connected neural networks (Haykin, 1994). We denote $h_{i}^{\ell}$, $\sigma_{i}^{\ell}(\cdot)$, $z_{i}^{\ell}$ as the input after linear transformation, activation function, and output of the $i$-th hidden unit at $\ell$-th layer, respectively. Then, we may formally define the computation within neurons as follows:
@@ 2.1. Activation and Monosemantic
$\forall_{\mathbf{x}}act(z_{i},\mathbf{x})=1,\mathbf{x}\in C_{j};\forall_{\mathbf{x}}act(z_{i},\mathbf{x})=0,\mathbf{x}\notin C_{j}$
However, features are human-defined and vary a lot. In the previous study, Gurnee et al. (2023) considered about 100 features, while Trenton et al. (2023) studied up to $10^{5}$ features to fully capture monosemanticity. Without unified feature datasets, it is also difficult to explicitly formalize the definition of “monosemantic”. More related works are discussed in Appendix B.
@@ 3.1. Metric for Monosemanticity
The proof is given in Appendix D.1. In the implementation, we introduce a forgetting mechanism during model updates, where the influence of previous samples should decay. For details, please refer to Algorithm 1 in Appendix E.2.
@@ 3.2.1. Naive Deactivation
Recall that given a monosemantic neuron $z$ and an input $\mathbf{x}$ that contains its exclusive feature, we aim to optimize the model in two aspects: (1) decrease the activation level of $z$ when given $\mathbf{x}$ (Figure 3(b)); and (2) reduce the reliance of output $\mathbf{o}$ on the activation of $z$ (Figure 3(c)).
@@
where subscript $\cdot_{ng}$ refers to a “no gradient” fixed-value scalar without trainable parameters, compared with $z\in\mathbf{z}=f_{1}(\mathbf{x})$ which is updatable. Two examples are given in Figure 3(e,f). Both solutions ensure that $f_{2}$ receives an updated $z^{\prime}$, in which the value of $\mathcal{V}(z^{\prime})=\bar{z}$ provides little information. As $z^{\prime}$ is valueless, $f_{2}$ has to adjust its parameter to utilize information from other neurons in $\mathbf{z}$ for good output $\mathbf{o}^{*}=f_{2}^{*}(\mathbf{z}^{\prime},\mathbf{x})$. Such a process achieves our second goal of reducing the dependence of output $\mathbf{o}$ on the activation of $z$ (Figure 3(l)).
However, neither of the two solutions can achieve the first goal (i.e., training $f_{1}$ to inhibit the activation of $z$ for a single feature). To be more specific, method (a) wastes all the gradient for $f_{1}$ in generating $z$, preventing the related parameters from being updated. Method (b) outputs $\mathcal{V}(z^{\prime})=\bar{z}$ by compensating for the gap. As the model obtains a good result with $z$, by calculating the gradient, $f_{1}$ will be updated to push its output value from $\bar{z}$ to $z$, expressed as $\bar{z}+\delta(z-\bar{z})$, where $\delta>0$ up to the learning rate (Figure 3(i)). Ironically, the actual output of $z$ is updated to $z+\delta(z-\bar{z})$, deviating from $\bar{z}$ by a factor of $1+\delta$:
@@
Therefore, these two simple solutions either contribute nothing to the deactivation of $z$ or even enhance its activation (Figure 3(g,k)).
@@ 3.2.2. Reversed Deactivation
To address the aforementioned problems, we propose our method called Reversed Deactivation (RD for short). Following the above-mentioned definitions, we replace the original $z$ with $z^{\prime}$ (Figure 3(f)):
@@
Similar to baselines, RD also ensures the second goal of decreasing the dependence of output $\mathbf{o}$ on the activation of $z$. To be more specific, $f_{2}$ receives a value $\mathcal{V}(z^{\prime})=\mathcal{V}({-z+(\overline{z}+z)_{ng}})=\overline{z}$, which provides little information and requires $f_{2}$ to learn the information from other neurons in $\mathbf{z}$ (Figure 3(h)).
Besides, RD can inhibit the activation level of $z$ when given $\mathbf{x}$. In short, after calculating the gradient, $f_{1}$ will update the trainable parameter to push its output value from $\bar{z}$ to $z$ (Figure 3(i)). Different from method (b), the gradient on $z$ is reversed. To achieve the same shifted value $\bar{z}+\delta(z-\bar{z})$ mentioned above, an insight into the update can be expressed as:
@@
which achieves deactivation (Figure 3(j)). The formal and detailed theory is presented in the following lemma.
Lemma 3.3.
Given a trained model $f$ with a continuous second derivative and a Lipschitz continuous gradient, where $f$ achieves a desired output $\mathbf{o}$ with minimal loss $\mathcal{L}(\mathbf{o})$, in which $\textbf{o}=f(\mathbf{x})=f_{2}(f_{1}(\mathbf{x}),\mathbf{x})=f_{2}(\mathbf{z},\mathbf{x})$ for input $\mathbf{x}$ based on its monosemantic neuron $z$ in layer $\mathbf{z}$, suppose that $\mathcal{L}(f_{2}(\cdot))$ monotonically increases with $|z^{\prime}-z|$ for any other value $z^{\prime}$ that replaces $z$. Then, with a sufficiently small learning rate $l$, by updating the model $f$ with gradient descent based on the neuron processed by the RD method, the activation of $z$ on input $\mathbf{x}$ can be inhibited.
The proof is given in Appendix D.2.
Additionally, we conduct validation experiments to compare the optimization outputs based on different methods in subsection 4.3. The results are consistent with the theory and the analysis of naive methods and RD, showing that RD is effective in inhibiting monosemanticity.
@@ 3.3. Flexible Plug-in Module
The details of MEmeL are displayed in Figure 4. We present a general framework of a neural network in Figure 4(a). The output $\mathbf{o}$ is derived from $\mathbf{x}$ based on the layers of neurons $\{\mathbf{z}\}$. Yellow arrows indicate the reliance of neuron layers on each other. Our method can be applied to any layer of neurons, such as $\mathbf{z}^{3}$ and $\mathbf{z}^{5}$ in Figure 4(b), where the original neurons are adjusted by our modules to inhibit monosemanticity.
Taking $\mathbf{z}^{5}$ as an example (Figure 4(c)), we first detect monosemantic neurons using our MS metric, as introduced in subsection 3.1. After identifying these neurons (colored red), we apply our Reverse Deactivation method, as described in subsection 3.2, to each of them.
@@
For lightweightness, neither of our two methods introduces any trainable parameters. Additionally, MEmeL focuses on presenting the idea of functionality induction. The methods we propose do not directly decrease monosemanticity during training, but instead induce the model to do so, supported by theoretical analysis. Once we have a well-trained model, the performance is expected to be robust even without induction. Therefore, during testing, the module can be directly removed without incurring any inference overhead. The results of the validation experiments also support our analysis of removing MEmeL during testing. (see Appendix F.1).
@@
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 68.2 80.2
MEmeL-Tune 84.8/83.9 71.7 91.2 93.7 55.7 86.6 89.0 68.2 80.5
The detailed algorithm is provided in the Appendix C.2. We also apply two simple tricks for implementation: late start and variance compensation, to avoid unstable value fluctuations during the cold start.
@@ 4. Experiments
In this section, we first introduce the basic experimental setup in subsection 4.1, then present the main empirical studies in subsection 4.2. We show a case study that the proposed Reverse Deactivation is more powerful in inhibition monosematicity than the naive methods in subsection 4.3. Furthermore, we discuss the potentials and limitations of MEmeL in subsection 4.4. Note the experiment about the impact of removing MEmeL during test is provided in Appendix F.1 due to the limited space.